A satellite can fail for many possible reasons, two of which are computer failure and engine failure. F or a given mission, it is known that: The probability of engine failure is 0.008. The probability of computer failure is 0.001. G iven engine failure, the probability of satellite failure is 0.98. G iven computer failure, the probability of satellite failure is 0.45. G iven any other component failure, the probability of satellite failure is zero. (a) D etermine the probability that a satellite fails. (b) D etermine the probability that a satellite fails and is due to engine failure. (c) Assume that engines in different satellites perform independently. G iven a satellite has failed as a result of engine failure, what is the probability that the same will happen to another satellite?


1
Expert's answer
2021-11-09T14:30:52-0500

Define the following events,

Let,

"C" be the event of a computer failure

"E" be the event of an engine failure

"S" be the event of a system failure

We have the following probabilities,

p(E)=0.008

p(C)=0.001

and the conditional probablities are given as,

p(S|E)=0.98

p(S|C)=0.45


a.

The probability that a satellite fails can be found using the law of total probability.

"p(S)=p(S|E)*p(E)+p(S|C)*p(C)"

"p(S)=(0.008*0.98)+(0.001*0.45)=0.00829"

Probability that the system fails is 0.00829


b.

 Probability that a satellite fails and is due to engine failure is given as "p(E\cap S)"and can be found using the formula below

"p(E\cap S)=p(S|E)*p(E)=0.008*0.98=0.00784"

Therefore, the probability that the system fails and is due to engine failure is 0.00784


c.

If engines in different satellites work independently then "p(E\cap S)" will be equal for each satellite therefore,

"p(E\cap S)=p(E| S)*p(S)=0.008*0.98=0,00784"

Given a satellite has failed as a result of engine failure, the probability that the same will happen to another satellite is 0.00784.